Integration Procedures(Itevrations by Parts)


∫ y √(y+1)dy


1
Expert's answer
2021-04-20T16:36:38-0400
"\int y\sqrt{(y+1)}dy=\int (y+1-1)\sqrt{(y+1)}dy"

"=\int (y+1)^{3/2}dy-\int \sqrt{(y+1)}dy"


"=\dfrac{2}{5} (y+1)^{5/2}-\dfrac{2}{3} (y+1)^{3/2}+C"

"=\dfrac{2}{15} (y+1)^{3/2}(3y+3-5)+C"

"=\dfrac{2}{15} (y+1)^{3/2}(3y-2)+C"